A 23.0 cm × 26.1 cm rectangular loop of wire carries 1.0 A of current clockwise around the loop. 23.0 cm 26.1 cm Part 1 out of 2 (a) Find the magnetic force on each side of the loop if the magnetic field is 2.0 T to the left (in the -x- direction). N in the (select) - direction top F bottom F left direction N in the (select) N in the (select) ? direction N in the (select) ? direction right
Answer #1

The force by the top and bottom wire is zero as they are parallel to the magnetic field.

Force in the left wire = ILB*sin (theta)

Fleft = 1*0.23*2 = 0.46 ( as per right hand rule, the direction of force is out of the page)

Fright = 1*0.23*2 = - 0.46 ( as per right hand rule, the direction of force is into of the page)

Note that the net force on the loop is zero.